[figure: 9.7 cm]

SUPERCONDUCTING AND MAGNETIC PHASES IN THE PEROVSKITE-LIKE CUPRATES: THE PROBLEM OF PHASE

SEPARATION

E.P. Khlybov^{1,2}, I.E. Kostyleva^{1}, T. Palewski^{2}, V. Nizhankovskii^{2}, K. Nenkov^{2,3}, J. Warchulska^{2}

1Vereshchagin Institute of High-Preassure Physics, Russian Academy of Sciences,

Troitsk, Moscow oblast, 142092 Russia 2International Laboratory of High Magnetic Fields and Low Temperatures, ul. Gajowicka 95, 53–421 Wroclaw, Poland 3Institut für Festkörper und Werkstofforschung Dresden, P.O. Box 270016, 01171 Dresden, Germany

Studies of magnetic and transport properties of the 1212 compounds MA_{2}LnCu_{2}O_{7−x} (M = Al or Tl; A = Sr or Ba; Ln = lanthanides or Y) revealed a number of interesting magnetic phenomena. Apart from high-temperature superconductivity, these compounds exibit spontaneous antiferromagnetism, field-induced antiferromagnetism or ferromagnetism. The different magnetic orders and superconductivity may be realized in the same compound as a result of the phase separation.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : T. Palewski

Address for correspondence : International Laboratory of High Magnetic Fields and Low Temperatures, ul. Gajowicka 95, 53–421 Wroclaw, Poland tel: 61-34 69(71); fax:61 27 21(71)

Email address : paltom@alpha.mlspmint.pan.wroc.pl